# Energy Rank Alignment: Using Preference Optimization to Search Chemical Space at Scale

### v1

@@ Abstract
  • Searching through chemical space is an exceptionally challenging problem because the number of possible molecules grows combinatorially with the number of atoms. Large, autoregressive models trained on databases of chemical compounds have yielded powerful generators, but we still lack robust strategies for generating molecules with desired properties. This molecular search problem closely resembles the “alignment” problem for large language models, though for many chemical tasks we have a specific and easily evaluable reward function. Here, we introduce an algorithm called energy rank alignment (ERA) that leverages an explicit reward function to produce a gradient-based objective that we use to optimize autoregressive policies. We show theoretically that this algorithm is closely related to proximal policy optimization (PPO) and direct preference optimization (DPO), but has a minimizer that converges to an ideal Gibbs-Boltzmann distribution with the reward playing the role of an energy function. Furthermore, this algorithm is highly scalable, does not require reinforcement learning, and performs well relative to DPO when the number of preference observations per pairing is small. We deploy this approach to align molecular transformers to generate molecules with externally specified properties and find that it does so robustly, searching through diverse parts of chemical space. While our focus here is on chemical search, we also obtain excellent results on an AI supervised task for LLM alignment, showing that the method is scalable and general.
@@ 1 Introduction
We formulate a generic alignment algorithm that we call Energy Rank Alignment or ERA that leverages an explicit reward function to guide autoregressive sampling while targeting specific properties or preferences. Unlike reward maximization in RL-based algorithms, the policy that minimizes our objective is designed to sample fluctuations around a maximal reward value to promote sample diversity. Our algorithm enables direct gradient-based optimization of a policy to match the ideal preference distribution and converges asymptotically to an optimal distribution with tuneable entropy and controllable regularization, which we show theoretically. The minimizers of our objective are closely related to the minimizer of PPO and DPO, but we have more direct control over the influence of the regularization relative to fluctuations around the maximum reward. In numerical experiments, we demonstrate that this algorithm successfully aligns a molecule transformer model to identify a highly diverse set of chemicals with properties favored by our choice of reward. Finally, we also show that we obtain competitive performance with ERA on benchmark LLM alignment tasks, but emphasize that the chemical applications are the main focus of this paper.
@@ 2 Energy rank alignment
and $\delta_{\pi}$ denotes the Fréchet derivative with respect to $\pi$. Assuming that $\pi_{0}$ has full support on $\mathcal{X}\times\mathcal{Y}$, the optimization converges asymptotically to stationary policy which satisfies
@@
Proximal Policy Optimization (PPO) optimizes an indirect, proximal objective to minimize an objective closely related to (2) (cf. Appendix B). Direct Preference Optimization (DPO) treats the negative reward function $U$ implicitly and directly maximizes the likelihood of $p(\boldsymbol{y}\succ\boldsymbol{y}^{\prime}|\boldsymbol{x})$. Our objectives differ from both approaches: like DPO, we directly optimize the policy using an explicit, gradient-based objective, but, in contrast, we use a reward function directly in our objective. The losses we build are thus amenable to both offline (samples from $\pi_{\rm ref}$) and online (samples from $\pi_{\boldsymbol{\theta}}$) policy alignment, as explained below. Choosing to optimize the objective online has been shown to have important consequences on performance [32], though we focus here on the setting where samples are drawn offline.
@@
instead of (9). One issue that arises with this scheme is that differentiation with respect to the parameters of the policy $\boldsymbol{\theta}$ because $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ are decoded into discrete tokens, an operation that is not differentiable. To remedy this, we importance sample with a reference policy
@@
3 Theoretical Analysis
@@
We prove Lemma 3.1 in Appendix B and use this simple lemma to prove the following result.
Proposition 3.2
@@
First, we verify that any probability measure $\pi_{g}(\boldsymbol{y}|\boldsymbol{x})\propto\exp(-\frac{\beta}{1+\gamma}\bigl{(}U(\boldsymbol{x},\boldsymbol{y})-\beta^{-1}\gamma\log\pi_{\rm ref}(\boldsymbol{y}|\boldsymbol{x})\bigr{)}+g(\boldsymbol{x}))$ minimizes the objective. Because $\mathcal{L}^{\mathrm{ERA}}$ is non-negative, it suffices to show that for all pairs $\boldsymbol{y},\boldsymbol{y}^{\prime}$, $D_{\rm KL}^{(\boldsymbol{y},\boldsymbol{y}^{\prime})}(p_{\gamma}|p_{\boldsymbol{\theta}})\equiv 0.$ This follows immediately from the cancellation in the preference probability $p_{\gamma}$ of $e^{g(\boldsymbol{x})}$ after factorization in (5). Now, suppose that $\pi(\boldsymbol{y}|\boldsymbol{x})\neq\exp{\left(-\frac{\beta}{1+\gamma}\bigl{(}U(\boldsymbol{x},\boldsymbol{y})-\beta^{-1}\gamma\log\pi_{\rm ref}(\boldsymbol{y}|\boldsymbol{x})\bigr{)}\right)}$ where we have taken $g(\boldsymbol{x})=0$ without loss of generality and $\pi:=\pi_{g}$. Assume that for all pairs $\boldsymbol{y},\boldsymbol{y}^{\prime}$, the divergence $D_{\rm KL}^{(\boldsymbol{y},\boldsymbol{y}^{\prime})}(p_{\gamma}|p_{\boldsymbol{\theta}})\equiv 0$ which is required of a minimizer. Equivalently, it must be the case that for all $\boldsymbol{y},\boldsymbol{y}^{\prime}$,

### v2

# Aligning Transformers with Continuous Feedback via Energy Rank Alignment

@@ Abstract
  • Searching through chemical space is an exceptionally challenging problem because the number of possible molecules grows combinatorially with the number of atoms. Large, autoregressive models trained on databases of chemical compounds have yielded powerful generators, but we still lack robust strategies for generating molecules with desired properties. This molecular search problem closely resembles the ”alignment” problem for large language models, though for many chemical tasks we have a specific and easily evaluable reward function. Here, we introduce an algorithm called energy rank alignment (ERA) that leverages an explicit reward function to produce a gradient-based objective that we use to optimize autoregressive policies. We show theoretically that this algorithm is closely related to proximal policy optimization (PPO) and direct preference optimization (DPO), but has a minimizer that converges to an ideal Gibbs-Boltzmann distribution with the reward playing the role of an energy function. Furthermore, this algorithm is highly scalable, does not require reinforcement learning, and performs well relative to DPO when the number of preference observations per pairing is small. We deploy this approach to align molecular transformers and protein language models to generate molecules and protein sequences, respectively, with externally specified properties and find that it does so robustly, searching through diverse parts of chemical space.
@@ 1 Introduction
We formulate a generic alignment algorithm that we call Energy Rank Alignment or ERA that leverages an explicit reward function to guide autoregressive sampling while targeting specific properties or preferences. Unlike reward maximization in RL-based algorithms, the policy that minimizes our objective is designed to sample fluctuations around a maximal reward value to promote sample diversity. Our algorithm enables direct gradient-based optimization of a policy to match the ideal preference distribution and converges asymptotically to an optimal distribution with tuneable entropy and controllable regularization, which we show theoretically. The minimizers of our objective are closely related to the minimizer of PPO and DPO, but we have more direct control over the influence of the regularization relative to fluctuations around the maximum reward. In numerical experiments, we demonstrate that this algorithm successfully aligns a molecule transformer model to identify a highly diverse set of chemicals with properties favored by our choice of reward. Finally, we demonstrate that ERA is able to align a protein language model to generate mutated protein sequences with desirable properties according to a computational reward model.
@@ 2 Energy rank alignment
and $\delta_{\pi}$ denotes the Fréchet derivative with respect to $\pi$. Assuming that $\pi_{0}$ has full support on $\mathcal{X}\times\mathcal{Y}$, the optimization converges asymptotically to a stationary policy which satisfies
@@
Proximal Policy Optimization (PPO) optimizes an indirect, proximal objective to minimize an objective closely related to (2) (cf. Appendix B, C). Direct Preference Optimization (DPO) treats the negative reward function $U$ implicitly and directly maximizes the likelihood of $p(\boldsymbol{y}\succ\boldsymbol{y}^{\prime}|\boldsymbol{x})$. Our objectives differ from both approaches: like DPO, we directly optimize the policy using an explicit, gradient-based objective, but, in contrast, we use a reward function directly in our objective. The losses we build are thus amenable to both offline (samples from $\pi_{\rm ref}$) and online (samples from $\pi_{\boldsymbol{\theta}}$) policy alignment, as explained below. Choosing to optimize the objective online has been shown to have important consequences on performance [38], though we focus here on the setting where samples are drawn offline.
@@
instead of (9). One issue that arises with this scheme is differentiation with respect to the parameters of the policy $\boldsymbol{\theta}$ because $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ are decoded into discrete tokens, an operation that is not differentiable. To remedy this, we importance sample with a reference policy
@@
3 Experiments
We test ERA on both chemical and language tasks to shed light on the following questions: 1) Can we use ERA to robustly fine-tune our model to generate samples according to a desired distribution? 2) What is the effect of changing the inverse-temperature $\beta$ during ERA? 3) Do we maintain sample diversity (and validity) without regularizing to remain close to a reference policy, and what is the effect of increased regularization? 4) Can we simultaneously target multiple properties with high fidelity, and how can we trade off between desired properties?
3.1 Generating small molecules with desired properties
We use a decoder-only representation for the molecular generator [4], where the generator has 2 layers, an embedding dimension of 512, a vocabulary of 324 tokens, and totals 3.5M parameters. Starting from a random initialization, we carry out pretraining on a dataset of 2.4M small molecules from the ChEMBL database [46] for 180 epochs. This version of the model is not conditioned on a prompt and generates a small molecule (represented as a SMILES string) given just a start-of-sequence token.
Central to ERA is, of course, access to a computable energy function. As a proof-of-concept, we first consider 5 different properties for which the corresponding energy function is easily evaluable: Quantitative Estimate of Drug-Likeness (QED) [6], Wildman-Crippen LogP (LogP) [43], Ring Count, Molar Refractivity (MR) [43], and Tanimoto Similarity [30] (Section 3.1.13.1.2). Briefly, LogP is a measure of the hydrophobicity of a molecule, MR is a measure of the polarizability of the molecule, and Tanimoto similarity is a measure of the similarity between two molecules (see Appendix E.2). We then investigate ERA in a more challenging context: generating small-molecules with high predicted bioactivity for two kinases (Section 3.1.3).
3.1.1 Unprompted molecular alignment on RDKit oracles
First, we independently target four different properties using ERA with an unprompted molecular generator (Fig. 2). Using the pretrained model as our reference policy, we generate a dataset $\mathcal{D}=\{\boldsymbol{y}_{1}^{(i)},\boldsymbol{y}_{2}^{(i)},U(\boldsymbol{y}_{1}^{(i)}),U(\boldsymbol{y}_{2}^{(i)})\}_{i=1}^{N}$ and carry out energy rank alignment on $\pi_{\boldsymbol{\theta}},$ where $\pi_{\boldsymbol{\theta}}$ is initialized using the weights of $\pi_{\rm ref}.$ Here, $\boldsymbol{y}_{1},\boldsymbol{y}_{2}\sim\pi_{\rm ref}$ and $\boldsymbol{y}$ and $U(\boldsymbol{y})$ denote the generated molecule and its corresponding energy, respectively. For MR, Ring Count, and LogP, we define the energy $U$ to be a harmonic potential centered at a target value. For QED, we define the energy to be the negative logarithm of QED and vary $\beta$ to assess its impact on alignment (see Tables S2S3). In Fig. 2, we see that we successfully shift the distribution to target means that are both greater and lower than the average value of MR, Ring Count, and LogP under the reference policy. Furthermore, in the alignment of QED, we observe the effect of changing $\beta$ on the learned policy; with increased $\beta,$ the learned policy concentrates around low-energy samples (i.e. near $\textrm{QED}=1$), and with lower $\beta,$ the learned policy samples a greater range of QED values, as expected. We note that for each of these four experiments, we did not regularize towards the reference policy (i.e. $\gamma=0$). Even so, we were able to maintain both sample diversity and maintain appreciable sample validity (see Fig. S8 and Table S4).
Many molecular design tasks require balancing multiple properties, and designing an objective for multi-property alignment is straightforward within the ERA framework. To demonstrate this, we generate molecules with both high QED and LogP using ERA with an energy function weighted by property-specific $\beta$: $U=\beta_{\rm QED}U_{\rm QED}+\beta_{\rm LogP}U_{\rm LogP}$ (see Tables S2S5 for details on the energy function). We carry out ERA with different pairs of $(\beta_{\rm QED},\ \beta_{\rm LogP})$ using the same procedure as above, and from Fig. 3, we see that we target multiple properties with varying fidelity by simply modulating the value of property-specific $\beta.$ Ultimately, increasing the $\beta$ for an individual property enables us to favor higher values of that property in multi-property alignment setting. In this case, we also do not regularize with the KL-divergence to the reference policy and again maintain sample diversity and validity (see Fig. S9 and Table S5)
3.1.2 Prompted molecular alignment on RDKit oracles
Inspired by the task of lead optimization in drug discovery efforts [21], we ask whether we can use ERA to train a molecular generator that can sample a molecule that is both similar to the prompt molecule and also exhibits some desired property. First, we fine-tune the pretrained molecular generator to enable prompted molecular generation (see Appendix E.3.2) and use this fine-tuned model as our reference policy for all prompted molecular alignment tasks. This reference policy disproportionately samples molecules that are identical (i.e. a Tanimoto similarity of 1.0) to the prompt molecule (see Fig. S10), so we carry out multi-property alignment on this reference policy to generate molecules that are similar—but not identical—to the prompt molecule and also have a high drug-likeness as measured by QED. Using ERA, we optimize the reference policy with a generated dataset $\mathcal{D}=\{(\boldsymbol{y}_{1}^{(i)},\boldsymbol{x}^{(i)}),(\boldsymbol{y}_{2}^{(i)},\boldsymbol{x}^{(i)}),U(\boldsymbol{y}_{1}^{(i)},\boldsymbol{x}^{(i)}),U(\boldsymbol{y}_{2}^{(i)},\boldsymbol{x}^{(i)})\}_{i=1}^{N}$, where we sample four molecules for each prompt molecule from the reference policy and consider all possible preference pairs for a total of six preference pairs per prompt molecule (see Appendix E.2 for full details on the energy used).
We observe that the per-prompt average QED under the optimized policy for a given prompt is higher than the corresponding average under the reference policy (Fig. S10). Furthermore, we see that we are able to sample a diverse set of molecules that are chemically similar to the prompt molecule, and also chemically valid (see Fig. S11, Table S6). We repeat the experiment with a related objective of generating molecules similar to the prompt molecule with a high LogP instead and again observe that we increase the per-prompt average LogP under the optimized policy relative to the reference policy without degrading sample diversity and validity. For both of these experiments, we required regularization to the reference policy ($\gamma>0$). With no regularization, the aligned generator would almost exclusively sample sequences that were chemically invalid ($<25\%$ chemical validity). Finally, we note that the increases in QED and LogP in Fig. S10 are smaller relative to the increases in Fig. 2 because the samples are now conditioned to remain proximal to the prompt molecule, which restricts the chemical space that can be explored.
3.1.3 Unprompted molecular alignment on protein-ligand docking oracles
[…]
mean score IntDiv mean score IntDiv
ERA 0.996 $\pm$ 0.000 0.219 $\pm$ 0.002 0.987 $\pm$ 0.001 0.264 $\pm$ 0.005
MolRL-MGPT 1.000 $\pm$ 0.000 0.362 $\pm$ 0.015 0.961 $\pm$ 0.010 0.372 $\pm$ 0.025
GFlowNet 0.649 $\pm$ 0.072 0.715 $\pm$ 0.104 0.437 $\pm$ 0.219 0.716 $\pm$ 0.145
GraphGA 0.919 $\pm$ 0.016 0.365 $\pm$ 0.024 0.875 $\pm$ 0.025 0.380 $\pm$ 0.015
JT-VAE 0.235 $\pm$ 0.083 0.770 $\pm$ 0.067 0.159 $\pm$ 0.040 0.781 $\pm$ 0.127
REINVENT 0.965 $\pm$ 0.011 0.308 $\pm$ 0.035 0.942 $\pm$ 0.019 0.368 $\pm$ 0.021
[…]
We additionally measure sample efficiency using the top-10 AUC metric [14], which is the area under the curve (AUC) of the mean property value of the top-10 performing molecules versus the number of oracle calls (see Fig. S13 and Table S7). We likewise only include novel, unique, and valid molecules in this analysis; any sampled molecule that is in ChemBL, that has already been sampled, or that is invalid is discarded and additionally does not count towards an oracle call as these are filtered out before oracle evaluation. We observe that we are able to generate novel and unique high-scoring molecules, with high sample efficiency especially in comparison to existing state-of-the-art methods. Ultimately, high sample efficiency is crucial in settings where evaluation is expensive, which will generally be true for most real-world chemical and biological tasks (e.g. wet-lab experiment). Finally, we also perform Glide Standard Precision [13] docking on the top-scoring molecules according to the oracles (score of 1.0) against their respective receptors. We observe that the diverse set of sampled molecules exhibit chemically plausable docked poses obtained from a physics-based docking approach (Fig. 4).
[…]
3.2 Directed evolution of proteins with ERA
Figure 5: Alignment of ESM3-1.4B with $\beta$=0, 0.1, 1.0, 10.0 and $\gamma$=0.001 on the task of maximizing EVmutation score. Positions 182, 183, 184, and 186 of the TrpB parent sequence were masked and ESM3-1.4B predicted amino acids at those sites. The distribution of the EVmutation scores for generated sequences shifts significantly as $\beta$ is increased.
We also consider the performance of ERA in a large-molecule setting, namely ML-guided directed evolution of proteins. Directed evolution campaigns aim to optimize a protein sequence toward some desired property of interest via iterative mutagenesis, library screening, and selection of best variants [41]. This has become a widely used methodology in protein engineering but comes with key limitations. The inherently iterative nature of directed evolution campaigns can lead to costly and time-consuming experimental campaigns, and meaningfully understanding the effects of protein mutations on protein activity can often be difficult. These challenges have led to the application of machine learning methods to more efficiently guide directed evolution campaigns [45] [36]. Given the success of ERA in guiding the optimization of small-molecules using a SMILES language representation, we examined whether ERA could be used to optimize large protein molecules using a protein language (i.e. primary sequence) representation.
There has been significant recent effort to design and train large protein language models (PLMs) [22, 17]. Furthermore, these models have demonstrated remarkable capabilities across a number of protein tasks [42, 36]. As such, we decided to use the state-of-the-art ESM3-1.4B [17] as our pretrained model, for which we carried out alignment using ERA. Despite the multimodal nature of ESM3, here, we only focus on generating primary-sequence-based representations of proteins.
We consider directed evolution of the $\beta$-subunit of tryptophansynthase (TrpB) from Thermotoga maritima, an enzyme that catalyzes tryptophan production [8]. Here, we seek to evolve the protein to increase its evolutionary fitness. In this work, we do not have access to experimental validation and so we evaluate the fitness of sequences using the computationally evaluable EVmutation score, an oracle that is predictive of a variant sequence’s performance relative to the parent sequence in its native function [18].
As in other directed evolution campaigns for the TrpB protein [44], we consider mutating four different sites to one of the 20 standard amino acids. We randomly sampled 512 mutated sequences, emulating a random mutagenesis experiment. Using ESM3-1.4B as our reference model, we carry out alignment using ERA with various $\beta=(0.1,1.0,10.0)$ and $\gamma=0.001$ and plot the results in Fig. 5 (see Appendix F for more details). We observe that with higher $\beta$, we are able to sample mutants with the highest possible EVmutation score in a single round of alignment. These results are promising for the application of ERA in directed evolution campaigns and future work will focus on the guidance of wet-lab directed evolution campaigns in conjunctions with multi-round, on-policy ERA.
4 Conclusions and Limitations
This paper introduces energy rank alignment, a simple and effective algorithm for policy optimization with an explicit reward model. We find that ERA is stable without extensive hyperparameter tuning, and sufficiently general to successfully align application-specific transformers for chemical search problems and protein language models. The algorithm exhibits strong performance with a variety of reward models, even ones with relatively weak signal. We analyze the minimizers of the ERA objective and find that they differ from the minimizers of popular policy alignment algorithms DPO and PPO in an important way: unlike PPO, the strength of regularization to the reference policy that we add is controlled by a parameter $\gamma$, while the entropy of the target distribution is independently tuned by a distinct parameter $\beta$. This means that we can avoid greedy policies by keeping $\beta$ small—amplifying fluctuations around the optimum of the reward model $(-U)$—while reducing the influence of the reference policy by taking $\gamma$ small. Our objective leads to easily interpretable sample-wise gradients which highlight the importance of a reward model relative to DPO in the sampled objective.
Limitations:
First, our approach requires a reward model, which can be difficult to train or design, especially for complex tasks. While we observed that ERA makes an appreciable impact even with weak supervision, this sort of proxy may not be available for more complex tasks. For example, optimizing small molecules for high binding affinity to a target protein would require expensive and noisy evaluations of a reward model, which likely limits the scope of molecular design to problems where the reward can be computed somewhat efficiently. A second limitation of our present work is that we do not train the molecular transformer to favor synthetic accessibility nor do we explicitly seek to obtain molecules that are easily synthesized experimentally. There are models that seek to evaluate synthesizability computationally that could be used in our rewards, which we plan to explore in future work [12].
Appendix A Code Availability
Appendix B Theoretical Analysis
@@
We prove Lemma B.1 in Appendix C and use this simple lemma to prove the following result.
Proposition B.2
@@
First, we verify that any probability measure $\pi_{g}(\boldsymbol{y}|\boldsymbol{x})\propto\exp(-\frac{\beta}{1+\gamma}\bigl{(}U(\boldsymbol{x},\boldsymbol{y})-\beta^{-1}\gamma\log\pi_{\rm ref}(\boldsymbol{y}|\boldsymbol{x})\bigr{)}+g(\boldsymbol{x}))$ minimizes the objective. Because $\mathcal{L}^{\mathrm{ERA}}$ is non-negative, it suffices to show that for all pairs $\boldsymbol{y},\boldsymbol{y}^{\prime}$, $D_{\rm KL}^{(\boldsymbol{y},\boldsymbol{y}^{\prime})}(p_{\gamma}|p_{\boldsymbol{\theta}})\equiv 0.$ This follows immediately from the cancellation in the preference probability $p_{\gamma}$ of $e^{g(\boldsymbol{x})}$ after factorization in (5).
Now, suppose that $\pi(\boldsymbol{y}|\boldsymbol{x})\neq\exp{\left(-\frac{\beta}{1+\gamma}\bigl{(}U(\boldsymbol{x},\boldsymbol{y})-\beta^{-1}\gamma\log\pi_{\rm ref}(\boldsymbol{y}|\boldsymbol{x})\bigr{)}\right)}$ where we have taken $g(\boldsymbol{x})=0$ without loss of generality and $\pi:=\pi_{g}$. Assume that for all pairs $\boldsymbol{y},\boldsymbol{y}^{\prime}$, the divergence $D_{\rm KL}^{(\boldsymbol{y},\boldsymbol{y}^{\prime})}(p_{\gamma}|p_{\boldsymbol{\theta}})\equiv 0$ which is required of a minimizer. Equivalently, it must be the case that for all $\boldsymbol{y},\boldsymbol{y}^{\prime}$,